You placed $2,227 in a savings account today that earns an annual interest rate of 4.75 percent, compounded semiannually. How much will you have in this account at the end of 9 years? Assume that all interest received at the end of the period is reinvested the next period.

Round the answer to two decimal places.
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=2,227*(1+0.0475/2)^(2*9)

=2,227*1.5257749

=$3397.90(Approx).